# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

### v4

@@ IV Conclusion
The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.3). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (13) and (16)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
@@
The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble and II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators.
@@
We wish to show that i) the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ of Eq. (32) is bounded from below by the Helmholtz free energy $F_{0}=-\text{k}_{\text{B}}T\log(Z)$ with $Z$ defined in Eq. (16), and ii) that the only minimizer of $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is $\mathcal{P}_{\text{eq}}$ of in Eq. (13). We will use Jensen’s inequality from convex analysis and probability theory to prove these two statements straightforwardly. This path is similar to that proposed by Friedli and Velenik in the case of classical lattice systems (see Lemma 6.74 in Ref. Friedli and Velenik, 2017) but here applied to the mixed QM/MM case. In order to be able to use Jensen’s inequality, we will first need to show that the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ is indeed a probability measure. While for $\boldsymbol{p}$ it is obvious, it requires a little more attention for $\hat{{\boldsymbol{\rho}}}$ as it consists formally in a operator. Then, $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ will be written as an average value on probability measures, and Jensen’s inequality and strict convexity will do the rest.
@@
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=$ $\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\big{(}\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}}\boldsymbol{p})\big{)}\}$ (A.12)
$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big{(}\epsilon({\bf{Q}},i)+\text{k}_{\text{B}}T\log(\mu({\bf{Q}},i))$
$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big{(}\epsilon({\mathbf{X}})+\text{k}_{\text{B}}T\log(\mu({\mathbf{X}}))$
$=$ $-\text{k}_{\text{B}}T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
$=$ $-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
@@
According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ at $\mathcal{P}_{\text{eq}}$ leads straightforwardly to
$F[\mu_{\text{eq}}]=$ $-\text{k}_{\text{B}}T\log(Z),$ (A.13)
which is of course the correct Helmholtz free energy. Then two questions remain: i) is $-\text{k}_{\text{B}}T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
@@
As the function $-\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer. The Jensen’s inequality can be stated as follows.
@@
As $-\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
@@
$-\text{k}_{\text{B}}T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
@@
$-\text{k}_{\text{B}}T\log(Z)\leq F[\mu],$ (A.19)
which guarantees the variational character. Eventually, the strict convexity of the $-\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into

### v5

@@ IV Conclusion
The assumptions of these derivations is a QM/MM partitioning of the system, together with models for both the MM part and the QM/MM interaction. We therefore do not discuss the QM/MM crossover nor how to go beyond the classical models. The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.3). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (14) and (17)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
@@
The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble, II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators, and III) an extension of the QM/MM functional to the semi-grand-canonical ensemble where the number of classical particles is allowed to vary.
@@
We wish to show that i) the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ of Eq. (35) is bounded from below by the Helmholtz free energy $F_{0}=-T\log(Z)$ with $Z$ defined in Eq. (17), and ii) that the only minimizer of $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is $\mathcal{P}_{\text{eq}}$ of in Eq. (14). We will use Jensen’s inequality from convex analysis and probability theory to prove these two statements straightforwardly. This path is similar to that proposed by Friedli and Velenik in the case of classical lattice systems (see Lemma 6.74 in Ref. Friedli and Velenik, 2017) but here applied to the mixed QM/MM case. In order to be able to use Jensen’s inequality, we will first need to show that the couple $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ is indeed a probability measure. While for $\boldsymbol{p}$ it is obvious, it requires a little more attention for $\hat{{\boldsymbol{\rho}}\,}$ as it consists formally in a operator. Then, $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ will be written as an average value on probability measures, and Jensen’s inequality and strict convexity will do the rest.
@@
$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=$ $\text{Tr}\big\{\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}\big(\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+T\log(\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p})\big)\}$ (A.12)
$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big(\epsilon({\bf{Q}},i)+T\log(\mu({\bf{Q}},i))$
$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big(\epsilon({\mathbf{X}})+T\log(\mu({\mathbf{X}}))$
$=$ $-T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
$=$ $-T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
@@
According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ at $\hat{\mathcal{P}}_{\text{eq}}$ leads straightforwardly to
$F[\mu_{\text{eq}}]=$ $-T\log(Z),$ (A.13)
which is of course the correct Helmholtz free energy. Then two questions remain: i) is $T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
@@
As the function $\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer. The Jensen’s inequality can be stated as follows.
@@
As $\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
@@
$T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
@@
$-T\log(Z)\leq F[\mu],$ (A.19)
which guarantees the variational character. Eventually, the strict convexity of the $\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into